Answer questions 17-19 below using the following information: A firm is producing in the short-run (with K fixed at 10 units) according to the following production function: F(L) = Q =10L -0.25L^2 . You also know that w=$15, r=$20 and the firm can sell its product for P=$7.50.

17. To minimize costs how much labor should the firm hire in the short run? a. L=16 b. L=30 c. L=20 d. None is correct

18. How much output Q is produced? a. Q=525 b. Q=300 c. Q=96 d. None is correct

19. How much is the company’s profit? a. Profit=$3287.5 b. Profit=$280 c. Profit=$1750 d. None is correct
Answer #1

Given

P=$7.50

w=$15

K=10

r=$20

17)

Q=10L-0.25L2

Marginal Product of labor=MPL=dQ/dL=10-0.50L

Marginal Revenue product=MRPL=P*(10-0.50L)=7.50*(10-0.50L)=75-3.75L

Set MRPL=wage rate for cost minimization

75-3.75L=15

3.75L=60

L=16

Correct option is

a. L=16

18)

Q=10L-0.25L2=10*16-0.25*16^2=96

Correct option is

c) Q=96

19)

Total Cost=rK+wL=20*10+15*16=$440

Total Revenue=P*Q=7.50*96=$720

Profit=Total Revenue-Total Cost=720-440=$280

Correct option is

b) $280